# High-dimensional Autoregressive Modeling for Time Series with Hierarchical Structures

### v1

@@ 5 Simulation studies
Figure 3: (a) Average MSEs under different action orders with varying intrinsic rank $r_{k}$. (b) Average rolling forecast MSEs for HOFM, HFM, and our proposed method under supervised factor modeling framework.
In the last experiment, we compare our proposed method with two hierarchical feature extraction methods: HFM (schmid1957development) and HOFM (yung1999relationship), under the framework of SFM. The data generating process follows (2.12) with the coefficient tensor $\mbox{$\mathscr{A}$}_{\mathrm{AR}}$ constructed by both HOFM and our proposed method. Details of the data generation and training procedure are relegated to Supplementary Material.
@@
To compare the one-step ahead forecast performance of HOFM, HFM, and our proposed method, we apply Algorithm 1 to estimate the coefficient tensor $\mbox{$\mathscr{A}$}_{\mathrm{AR}}$ for our model, and use the alternating least squares algorithm to estimate the coefficient tensor for HFM and HOFM. Figure 3(b) presents the average rolling forecast MSEs over 50 replications. We observe that our method consistently outperforms both HFM and HOFM, regardless of whether the data are generated by HOFM or our method. This demonstrating the efficiency of our model in capturing hierarchical structures with shared components within each mode in loading matrices for feature extractions.
6 Real data analysis
In this section, we demonstrate the application of our proposed model to the International Personality Item Pool (IPIP) NEO-120 personality inventory (johnson2014measuring), using data collected from June 1, 2015 to April 30, 2025. The dataset is available in the Hugging Face Datasets repository: https://huggingface.co/datasets/ecorbari/IPIP120-SCORES. It includes scores for the five major personality traits (Neuroticism, Extraversion, Conscientiousness, Agreeableness, and Openness to Experience), assessed by 24 questions. Participants are categorized by gender (two groups) and age (six groups: 10-19, 20-29, 30-39, 40-49, 50-59, and 60+), resulting in 12 demographic groups. To analyze trends in personality traits over time, we compute monthly average scores for each question (recorded on a 1-5 Likert scale) within each demographic group, forming a tensor time series $\{\mbox{$\mathscr{Y}$}_{t}\in\mathbb{R}^{12\times 5\times 24}\}_{t=1}^{T}$ with $T=115$ observations. There is one missing entry in the dataset, which we impute using spline interpolation. For preprocessing, the data are first detrended by first-order differencing, and then standardized to have zero means.
The tensor autoregression model (2.12) is employed with each coefficient tensor $\mbox{$\mathscr{A}$}_{l}\in\mathbb{R}^{12\times 5\times 24\times 12\times 5\times 24}$. Observations are divided into a training set with the first 103 months and a testing set with the last 12 months. To enable a feasible estimation, we consider our proposed model, SFM based on CP and TT decompositions, and HOFM under SFM frameworks for feature extraction with the lag order fixed at $L=1$ by jointly consider the PACFs of all entries in data. We omit the comparison with Tucker decomposition and HFM based on SFM frameworks due to thier large number of parameters in the coefficient matrix, which makes them infeasible for this data analysis. For our proposed model, from prior knowledge of the data, we set the candidate action order sets as $\mathcal{P}_{x}^{(0)}=\mathcal{P}_{y}^{(0)}=\{(3,2,1),(1,3,2)\}$, i.e., the features are extracted by the order of questions, personalities, and demographic groups, or the order of demographic groups, questions, and personalities. We then apply Algorithm 2 to get the selected intrinsic ranks $\boldsymbol{r}_{1}=(1,1,2)$, $\boldsymbol{r}_{2}=(9,1,2)$, $\boldsymbol{s}_{1}=(3,1,2)$, and $\boldsymbol{s}_{2}=(3,2,2)$. For the two tensor decomposition based SFMs, we adopt the ALS methods in zhou2013tensor for CP-based tensor autoregression, and wang2022high for Tucker-based tensor autoregression. For TT-based SFM, we follow si2024efficient and qin2025computational to perform Riemannian gradient descent for estimation. The intrinsic ranks for these methods are selected by their respective methods.
We next evaluate the rolling forecast performance on the testing set of these four methods, i.e., fit the model to historical data up to month $t$ to forecast the data at month $t+1$ for $t=103,\ldots,114$. Specifically, we use mean squared forecast errors (MSFE) and mean absolute forecast errors (MAFE) to assess the one-step ahead forecast accuracy. Table 2 summarizes the results, where our proposed method outperforms all competing methods in both MSFE and MAFE metrics. This result supports the existence of hierarchical structures in the IPIP NEO-120 data, which can be effectively captured by our proposed model. Moreover, the superior performance of our method over TT decomposition based SFM highlights the advantage of our model by considering multiple action orders for general hierarchical data. And the better performance compared to HOFM based SFM demonstrates the benefit of improved efficiency in feature extraction by sharing components within each mode in loading matrices. In conclusion, our proposed model provides an efficient and interpretable approach for modeling tensor time series data with hierarchical structures.
Table 2: Mean squared forecast errors (MSFE) and mean absolute forecast errors (MAFE) of our model and three competing methods on the testing set of the IPIP NEO-120 data. The smallest numbers in each row are in bold.
Ours CP $T$ HOFM
MSFE 138.78 142.28 145.73 142.15
MSAE 279.28 280.63 283.43 280.58
Finally, we zoom in on the estimated coefficient tensor from our proposed model to interpret the hierarchical structures in the IPIP NEO-120 data and the importance of different action orders. We discuss the component matrices involved in feature extraction for responses under action order $(1,3,2)$ as an example. Figure 4(a) presents the heatmap of the component matrix $\boldsymbol{H}_{1,1}^{\top}$, which correspond to the first step of feature extraction, summarizing the information across the 24 questions for each personality trait. In addition, Figure 4(b) displays $\boldsymbol{H}_{2,1}^{\top}$ that represents the combined effects of personality traits and questions by a single extracted feature. Lastly, Figure 4(c) illustrates $\boldsymbol{H}_{3,1}^{\top}$, capturing the interactions between demographic groups and the combined features from questions and personality traits. The response tensor $\mbox{$\mathscr{Y}$}_{t}$ is then summarized into two factors as its low-dimensional representation. Component matrices with the second action order $(1,3,2)$ for responses and those for predictors with both action orders can be found in Supplementary Material, which have similar interpretations as those for predictors. Finally, Figure 4(d) presents the heatmap of $\boldsymbol{\Theta}$, the estimated coefficient matrix that captures the relationships between the extracted factors from predictors and responses by two action orders. We observe that features extracted by action order $(3,2,1)$ have a more significant impact on extracted features from responses under both action orders compared to those extracted by action order $(1,3,2)$, as indicated by the larger absolute values in the corresponding blocks of $\boldsymbol{\Theta}$.
Figure 4: Heatmaps of (a) $\boldsymbol{H}_{1,1}^{\top}$, (b) $\boldsymbol{H}_{2,1}^{\top}$, (c) $\boldsymbol{H}_{3,1}^{\top}$, and (d) $\boldsymbol{\Theta}$.
7 Conclusion and discussion
In this paper, we introduce a novel autoregressive model under SFM framework for tensor time series data that effectively captures hierarchical structures through the incorporation of multiple action orders in the construction of the coefficient tensor. We employ a structure of shared components within each mode in loading matrices for efficient feature extraction. An alternating least squares algorithm is developed for parameter estimation, and a boosting-based method is proposed for hyperparameter selection, which aims to reveal the underlying hierarchical structures present in the data. Theoretical analyses demonstrate the consistency of our estimator under sub-Gaussian noise with misspecification of action orders. Simulation studies and real data analysis illustrate the superior performance of our method compared to existing approaches.
For future research, we may extend the proposed model to accommodate more complex time series structures, such as moving average (MA) or ARMA models. We can also develop robust estimation techniques for handling heavy-tailed noise distributions, so that our methods can be deployed to a wider range of real-world applications. Additionally, investigating statistical inference procedures for the selected action orders and intrinsic ranks would provide valuable insights into the interpretability and reliability of the model in practical settings.

### v2

@@ Appendix E Setups and Results of the Third Simulation Studies in Section 5
Figure 5: Average rolling forecast MSEs for HOFM, HFM, and our proposed method under our new supervised modeling framework.
@@
Algorithm 1 is employed to estimate our models, while an alternating least squares algorithm is used for models with HFM and HOFM being feature extracting tools. One-step ahead forecasting is used for the evaluation, and Figure 5 gives the MSEs, averaged over 500 replications. Our model consistently outperforms both HFM and HOFM, regardless of whether the data are generated by HOFM or our method. It perhaps is due to the fact that both HFM and HOFM involve large number of parameters. As a result, the efficiency of our model in handling hierarchical structures can be well demonstrated by capturing shared components within each mode under high-dimensional settings.
Appendix F Additional Analysis for the Personality‑120 Data
We first offer an interpretation of Item 58 for the first action order $\alpha_{1}=(3,2,1)$ as introduced in Section 6, following the way Item 46 is analyzed in the main text. For elderly females, the important questions include Question 3 and 9 of Neuroticism trait, as well as Question 4, 7, and 13 of Openness trait. These questions are summarized as follows: “Am afraid of many things” and “Often feel blue” exhibit positive loadings, signifying a tendency of anxiety and depression, while “Like to get lost in thought,” “Prefer variety to routine,” and “Do not like poetry” exhibit negative loadings, signifying a loss of imagination and adventurousness. These questions, despite being scattered across different traits, can be collectively interpreted as a concern of control and security among elderly females.
Figure 6: Heatmaps of obliquely rotated estimated component matrices and the coefficient matrix: (a) $\widehat{\boldsymbol{G}}_{1,2}^{\top}$, (b) $\widehat{\boldsymbol{G}}_{2,2}^{\top}$, (c) $\widehat{\boldsymbol{G}}_{3,2}^{\top}$. In (c), A, C, E, N, and O stand for Agreeableness, Conscientiousness, Extraversion, Neuroticism, and Openness, respectively.
For the second action order $\alpha_{2}=(1,3,2)$, we present the obliquely rotated component matrices for both predictors and responses to illustrate the interpretability of our proposed model. Figure 6 shows the heatmaps of the estimated component matrices $\widehat{\boldsymbol{G}}_{1,2}^{\top}$, $\widehat{\boldsymbol{G}}_{2,2}^{\top}$, and $\widehat{\boldsymbol{G}}_{3,2}^{\top}$, whose interpretation follows the same logic as Section 6. In Figure 6(a), Group 1–6 represent males and Group 7–12 females. Within each group, a larger index indicates a higher age, and therefore the elderly tend to have a different attitude than the younger, especially for females.